What is the domain of C{3te3 -- 2e2+ + sin(2t)}(s)? Os> -3 Os> -2 Os > 2 Os > 0 Os > -1 None
Answer #1

h[ 3 1 ² 3 4 - 2 et + sin(24)} = 32 (E² e 3t 1-2h [ ezt ] + h [Sin (2+)] = 3.21 3 - 2 1 - 2 + 3 2 2 44 s=3 and s=2 are not in